(23x+14)Squared=25
I took the square root of both sides and brought the right side of the equation over in order to rationalize the equation. I'm left with
23x-9=0
I need to solve for "x" am i on the right path.

Answer by stanbon(75887)

(23x+14)Squared=25
-----------
(23x+14)^2 - 25 = 0
Factor:
(23x+14-5)(23x+14+5) = 0
(23x+9)(23x+19) = 0
x = -9/23 or x = -19/23
=============================
Cheers,
Stan H.